The maximum current that flows through a fuse wire, before it blows out, varies with its radius 'r' as -


(a) r3/2  


(b) r


(c) r2/3


(d) r1/2


 


Please answer in detail

Grade:12

1 Answers

Najbul Gazi
3 years ago
Option -(a) r^(3/2) 
  The heat produce by the fuse wire= i^2R= i^2* row*L/(π*r^2) 
Now, the rate of loss of heat from the fuse wire by radiation= Emissivity* Area * Excess temp. Over surrounding=e*2πrL*theta
 
Before the fuse will be blown out, 
   The heat produce in fuse wire = rate of loss of heat from the fuse
So, i^2*row*L/(πr^2) =e*2πrL*theta
  i^2 = [e*2πrL*theta*πr^2]/[row*L]
 Clearly, i^2 proportion r^3
And i proportion to r^(3/2)